OpenStudy (anonymous):

what is the Quadratic equation if the given Vertex is (0,-2), and the given points within the parabola is (2,0), and (-2,0)? please help!

Directrix (directrix):

If h and k are the coordinates of the vertex, the quadratic function in vertex form corresponding to this parabola is given by f(x) = a (x - h)^ 2 + k (h,k) = (0,-2) coordinates of vertex f(x) = a ( x - 0)^2 + (-2) y = a(x^2) - 2 (2,0) is a point on the parabola 0 = a(2^2) -2 2 = 4a a = 2/4 = 1/2 Equation: y = 1/2 (x^2) - 2

Directrix (directrix):

Click on link to see graph: http://www.wolframalpha.com/input/?i=y+%3D+1%2F2+%28x%5E2%29+-2+

Directrix (directrix):

OK?

OpenStudy (anonymous):

thanks! ^_^

Directrix (directrix):

Any time - thanks for the medal. :)